A discussion of Aristotle’s De Anima

St John’s College alumni unofficial email list

begun 2004.2.8

1 (L. A. Salas). Is anyone interested in starting an online discussion on Aristotle’s De Anima?

I would also be interested in discussing the Iliad or Hesiod’s Theogony and Works and Days.

A final suggestion from Austin would be Plato’s Parmenides, on which I’m currently having a seminar and which is ridiculously weird and insane about halfway through the dialogue. It also has a very rich dramatic setting.

2 (Mr Salas). I would be willing to ask the first question. Before we start, though, I would like to see how many other listmembers are interested in De Anima and lay down some parameters for discussion. For one thing, do we want to begin with chapter Α and work through page by page until Γ? Should we coordinate editions and such? The easiest thing for non-Greek-readers would probably be to use the Barnes collected works (they’re good and common enough that the local library should have a set if one doesn’t want to purchase a set). We probably want to talk about commentaries and such. Depending on the interest in De Anima, I can post a general bibliography on it (including commentaries, ancient and modern, and additional resources) if that’s desirable. Also, since commentaries already begin to blur the line between primary and secondary texts, to what extent, if any, do we want to include secondary literature?

3 (S. Thomas). As Mr Salas knows, I am happy to read such things very slowly. In the case of the writings that have come down to us as writings of Aristotle, we generally don’t know precisely how and by whom the words were written, neither do we know the purpose of the texts. Accordingly, it may be less important to start at the beginning and go to the end, as it is entirely possible that the text we have known as De Anima was not conceived as a whole. Nevertheless, as it has been decades since I’ve looked at the work, I would say that we should start at the beginning and go to the end, bearing in mind that these very concepts may be misleading with this particular ‘book’.

2004.5.7 edition  De Anima discussion  (3)

I would say that coordinating English editions has its own set of drawbacks, as the clash of translations is sometimes quite illuminating. I also would resist any attempt to eschew commentaries where helpful, bearing in mind that a commentary lacks the authority of the text itself. But in a written seminar, there’s no reason why W.D. Ross or the like can’t serve as a part of the written seminar as well as the likes of me. And I would welcome a bibliography.

When do we begin?

4 (H. Peterson). All of the above sound good to me—especially Hesiod, which I’ve always meant to read. But I’m fine with starting with De Anima.

On translations: I agree that it’s helpful to use an edition with the same numbering scheme and also that different translations can illuminate one another. This being an imperfect world, we usually can’t have both at once, so I’d vote for the latter course. (Though a discussion of which translations are best is always nice.)

I’m on digest, so I don’t know whether anyone has posted this info yet, but here’s the only online translation I’ve been able to find:

J. A. Smith translation (three HTML pages, with Bekker page numbers) http:

//psychclassics.yorku.ca/Aristotle/De-anima/

Nothing at Perseus, alas. Anyone know anything about this translation? (Gets ready to cover eyes.)

Mr Thomas wrote: ‘As Mr Salas knows, I am happy to read such things very slowly.’ I don’t know whether you meant your remark that way, sir, but a hear-hear to our not going too quickly. I have to admit that the swiftness of our discussion about this—I’m offline for twenty-four hours, and already the opening question is about to be posted—is a bit too quick for me. Can we have some accommodation for those of us who aren’t online every day?

5 (Mr Salas). No worries, there’s no opening question about to spring forth. I don’t know what the others think about appropriate lag time but I think that we should all contribute at our own pace. The conversation should be, as our fancy t-shirts say, asynchronous. I doubt that we’ll all be able to coordinate our schedules. Listmembers more experienced in this sort of thing will let us know where the shoals are, though.

Although I would be interested in Hesiod, it looks like the De Animas have it so far. As Mr Thomas mentioned in an earlier post, most editions will print Bekker paragraphs. Bekker numbers make references fairly easy. The Bekker numbers refer to the Greek, though. So, we shouldn’t be surprised if our translations don’t match up perfectly. I’ll make a post on the translations, commentaries, Greek editions, etc… that I know, separately.

6 (Mr Salas). I’ll divide this post into editions of the text, then commentaries (philosophical and grammatical).

Greek and Greek/English:

The standard Greek edition can be found in the Oxford Classical Texts series (ed. W.D. Ross). [somewhere in the $30 range, I think]

The Loeb edition, published by Harvard University Press, contains Greek and English facing pages (trans. Hett). [$20—and in most libraries] The Greek text is inferior to the OCT text and the apparatus criticus is non-existent. However, this edition is probably the best for a reading in English with occasional reference to the Greek.

We may need to make reference to the Parva Naturalia also. The Loeb edition of De Anima contains PN listed under Minor Works—or so I think but it’s in there at any rate. I don’t know if there is an OCT of the Parva Naturalia but there is a good Greek text in the W.D. Ross commentary I will mention below.

English Translations

The Oxford Clarendon series has a great translation (Hamlyn) but it only covers books 2 & 3 [(with selected passages from book 1)]. It also contains a superb commentary, mostly philosophical.

Barnes’ complete works that I already mentioned are great and include all of Aristotle’s extant works, but are admittedly a little expensive [around $50].

Commentaries

The standard Greek commentary on De Anima is by W.D. Ross. It exists in many university libraries and as a reprint through Postscript Books/Sandpiper Press. If you want to wrestle with the Greek, this commentary is probably a must. It also has good philosophical commentary.

The commentary for the Parva Naturalia is also by Ross, if I’m not mistaken, and also rocks.

Both of these commentaries come with the Greek text included but with no translations.

That’s all I can think of right now. The Loeb is probably one of the best places to go, since it has both De Anima and the Parva Naturalia in it. It is also relatively inexpensive and allows the reader to refer to the Greek.

7 (Mr Thomas). Here’s what I’d propose: let’s give ourselves until the end of February to get texts and do whatever else. (In my case, I’m thinking I’ll read the Barnes English version quickly, and the start on the Greek in Ross’s commentary, but that’s just me.)

Then, starting in March we can begin discussing in small chunks, taking whatever time we need until we are done. Perhaps one of us can try to put the seminar into a continuous form, as Mr Pierce did with the aborted Kant discussion.

8 (Mr Salas). From the Aristotle bibliography of the joint program in ancient philosophy at the University of Texas, Austin:

Aristotle: Texts & general studies

OCT: standard modern editions for most works (esp. those ed. by W.D. Ross); some major omissions supplied by Teubner and Budé; most also in Loeb collection (beware most translations). Also Ross’ editiones maiores (with studies of mss. and English commentary) of Analytics, Physics, De Anima, Parva Naturalia, and Metaphysics.

I. Bekker, ed. Aristotelis Opera (l83l): Bekker’s pagination is standard for subsequent editions and citations.

H. Bonitz, ed. Index Aristotelicus (1870) = vol. 5 suppl. to Bekker: incomplete but useful guide to terminology.

O. Gigon, ed. Aristotelis Opera, Vol. III: Librorum Deperditorum Fragmenta (1987) = new vol. 3 of Bekker.

V. Rose, ed. Fragmenta (1886): standard but incomplete collection.

W.D. Ross, Fragmenta Selecta (1955): widely cited selection (lacks apparatus).

J. Barnes, ed., The Complete Works of Aristotle (1984): revised Oxford translation in 2 vols.

Clarendon Aristotle: exacting translations with philosophical commentary: J.L. Ackrill, Categories and De Interpretatione; J. Barnes, Posterior Analytics; R. Williams, Topics 1 & 8; W. Charlton, Physics 1–2; E. Hussey, Physics 3–4; C. Williams, On Generation and Corruption; D.W. Hamlyn, De Anima 2–3; D. Balme, Parts of Animals 1 and Generation of Animals 1; C. Kirwan, Metaphysics Γ�Ε; D. Bostock, Metaphysics Z & H; J. Annas, Metaphysics M–N; M. Pakaluk, Nicomachean Ethics 8–9; M.J. Woods, Eudemian Ethics 1–2 & 8; T. Saunders, Politics 1–2; R. Robinson, Politics 3–4; D. Keyt, Politics 5–6; R. Kraut, Politics 7–8.

J.L. Ackrill, Aristotle the Philosopher (1981): incisive introduction to his methods and ideas on central topics.

I. Düring, Aristotle in the Ancient Biographical Tradition (1957): collection of testimonia on life and activity.

W. Jaeger, Aristotle: Fundamentals of the History of His Development (1923; tr. 1948): seminal but flawed study.

I. Düring, Aristoteles (1966): comprehensive study in German; good bibliography by topics.

H. Cherniss, Aristotle’s Criticism of Plato and the Academy (l944): classic but severe critique of Aristotle.

T. Irwin, Aristotle’s First Principles (1989): comprehensive and complex study; good bibliography.

G.E.R. Lloyd, Aristotle: The Growth and Structure of His Thought (1968): readable survey.

J. Lear, Aristotle: The Desire to Understand (1988): stimulating interpretive survey.

W.D. Ross, Aristotle: A Complete Exposition of His Work & Thought (1930): comprehensive and judicious.

Aristotle: Collections of Articles

J. Barnes, M. Schofield, R. Sorabji, eds. Articles on Aristotle, 4 vols. (l975–79): classic papers on all areas of Aristotle’s thought; extensive but dated bibliographies arranged by topic.

J. Barnes, ed. Cambridge Companion to Aristotle (1995): concise topical studies, excellent recent bibliography.

J.M.E. Moravcsik, ed. Aristotle (1967): classic papers on some central issues, from analytic perspective.

A.O. Rorty, ed. Essays on Aristotle’s Ethics (l980): seminal papers.

A. Gotthelf and J. Lennox, eds. Philosophical Issues in Aristotle’s Biology (1987): influential papers.

A.O. Rorty and M. Nussbaum, eds. Essays on Aristotle’s De Anima, (1992): influential papers.

L. Judson, ed. Aristotle’s Physics (1991): influential papers.

A.O. Rorty, ed. Essays on Aristotle’s Rhetoric (1996): diverse philosophical perspectives.

Symposium Aristotelicum: papers from influential triennial conference: I. Düring & G. E. L. Owen, eds. Aristotle and Plato in the Mid Fourth Century (1960); S. Mansion, ed. Aristote et les problèmes de méthode (1961); G.E.L. Owen, ed. Aristotle on Dialectic: The Topics (1968); I. Düring, Naturphilosophie bei Aristoteles u. Theophrast (1969); P. Moraux & D. Harlfinger, eds. Untersuchungen zur Eudemischen Ethik (1971); P. Aubenque, ed. Études sur la Métaphysique d’Aristote (1979); G.E.R. Lloyd & G.E.L. Owen, eds. Aristotle on Mind and the Senses (1978); E. Berti, ed. Aristotle on Science: The Posterior Analytics (1981); P. Moraux & J. Wiesner, ed. Zweifelhaftes im Corpus Aristotelicum (1983); A. Graeser, ed. Mathematics und Metaphysik bei Aristoteles (1987); G. Patzig, ed. Aristoteles Politik (1990); D. Furley & A. Nehamas, ed. Aristotle’s Rhetoric (1994); M. Frede & D. Charles, ed. Aristotle’s ’Metaphysics’ Lambda (2000); ongoing.

Greek Commentators on Aristotle

Commentaria in Aristotelem Graeca (a.k.a. CAG) 23 vols. & 3 suppl. (l882– l909): most from 2d to 6th C a.d. Generally most helpful or instructive are those by Alexander of Aphrodisias, Simplicius, and John Philoponus; many have been influential in modern study of ancient philosophy, both for careful interpretations of Aristotle’s texts and for rich testimony on other ancient philosophers.

Large parts of many of these commentaries are being published in scholarly English translations by various hands under the general ed. of Richard Sorabji; about 25 volumes to date.

9 (Mr Thomas). In addition to the translation by Joe Sachs (Aristotle’s on the Soul and on Memory and Recollection, Green Lion Press, 2001), there are translations by Hippocrates G. Apostle, the Loeb by one W.S. Hett and a translation in the 1907 edition of the work by R.D. Hicks. (This edition was based on an earlier one, and it is not clear to me what the provenance of the translation is.)

To avoid confusion: W.D. Ross is also sometimes referred to as Sir David Ross. They are (or rather were) the same person, however.

10 (D. Pierce). Mr Thomas hoped somebody could mark up the coming Aristotle discussion. It would be nice. If I end up being interested in the discussion, I may do the job—but I shall grumble if people do not edit their emails carefully before sending. On the other hand, making the emails into a single document need not be done by a single person; there are ways to share the burden. Again, one way to share the burden is for all email-writers to follow certain typographical conventions. Beyond that, well… when I save emails in a folder called ‘jlist/Aristotle’ for example, they are really being added to the end of a single file. When I make this into a tex file, I have to delete all of the headers from the emails (while making a note of the writers). Anybody can do this—and this yields a file that can be printed out and be studied, though it may not look all that nice.

11 (Mr Salas). Dear soul brothers and sisters, I’m very sorry to have been so long in posing an opening question, this past month has been brutal in Austin academics.

Let me throw out two questions to start out with, if that’s OK. The first question I have no opinion on, whereas the second niggles at me.

Question one: What does Aristotle mean in 402a by claiming that knowledge of the soul is highest in virtue of its accuracy (�κρ�βεια)?

Prolegomena to the second question: Aristotle begins De Anima by describing the need for a methodology in studying the soul. In order to find the appropriate methodology, he considers three facets of the soul 402a7ff, nature (φ�σις), essence or being (ο�σ�α), and its attributes (�σα συμβ�βηκε). Regarding these attributes, Aristotle makes a positive claim, that some of these attributes are peculiar to the soul (�δια) and some belong to animals also, on account of the soul.

Second question: What do we make of these attributes, which Hett unpacks as ⟨essential⟩ attributes of the soul (does anyone see a textual reason to translate �σα συμβ�βηκε as essential attributes?), and what do we make of Aristotle’s claim, en passant, that there are some attributes peculiar to the soul and some that the animal shares? Does Aristotle commit himself to anything by claiming that some attributes are peculiar to the soul?

12 (Mr Pierce). Right now the only text I have is the translation at http://

classics.mit.edu/Aristotle/soul.1.i.html, which doesn’t give the Bekker

pages; but I think in your second question you allude to the fifth paragraph there.

In any case, Aristotle there seems to suggest that the soul could be like straightness. Straightness does not touch a bronze sphere at a point; a straight object does. Similarly, it seems, a soul does not have an emotion; an animated (besouled) body has an emotion.

But the analogy does not seem very apt. A plane may touch a sphere at a point; but a bronze sphere touches a flat board in a little patch, and will probably make an indentation in the board if the sphere is on top.

Whereas, an emotion such as anger requires a body, because there is no anger in the abstract; there is only this or that instance of anger, which cannot be separated from telltale signs like raised blood pressure.

There is a mathematics of spheres and planes; there does not seem to be a mathematics of ‘passion, gentleness, fear, pity, courage, joy, loving, and hating’.

Okay, so if not emotions, what attributes of an animal are properly attributes just of the soul? Have I overlooked examples?

13 (Mr Salas). Dear Mr Pierce, The second question refers to a line in the first paragraph of your text (it’s taken up in greater detail in the fifth). Your text has the line

Our aim is to grasp and understand, first its essential nature, and secondly its properties; of these some are taught to be affections proper to the soul itself, while others are considered to attach to the animal owing to the presence within it of soul.

The translation is a little blurry in the first clause. The Greek says that the nature (φ�σις), the definition or essence (ο�σ�α), and then the attributes (�σα συμβ�βηκε) of the soul are the objects of the inquiry.

Since some attributes, according to Aristotle, are peculiar to the soul and some are imbued by the soul onto [sic] the body. I think that, in your example, straightness may be argued to be a property of the flat board quâ plane figure. This instance of attribution would fall into the category of an attribute that the body acquires in virtue of something else, in this case the nature of the plane figure. I don’t know what a genuinely peculiar (�δια) property of the soul might be for Aristotle but part of the motivation for my question lay in this puzzling ‘exclusive quality.’ What a property peculiar to the soul may be is obscure to me but it’s curious enough that I wonder where Aristotle is going with his distinction.

You had mentioned also,

Whereas, an emotion such as anger requires a body, because there is no anger in the abstract; there is only this or that instance of anger, which cannot be separated from telltale signs like raised blood pressure.

I think that this question’s big. I don’t know what the etiquette is for our discussion but I would suggest that we hold off on this for a little bit.

14 (Mr Salas). I’m not sure whether this caveat has cropped up yet. If it hasn’t, it’s important to keep in mind that translating ψυχ� as soul is dangerous. I don’t know that it’s avoidable but the term soul brings with it baggage of which we may want to be cautious.

15 (B. Porter). I’m wondering why there has been so little response to Mr Salas’s opening questions. Here’s my effort to stir the pot, with hopes of attracting appetites.

I, too, wonder what Aristotle could possibly have been thinking when he starts out asserting that the science of the [ψυχ� or] psyche (dare we call it ‘psychology’?) is to be respected because of its accuracy (or ‘exactness’ in my translation). Is he setting expectations, that we will come out the other end of the treatise with an exact and accurate definition of what the soul is? In our academic world, math and physics do get a lot of respect for being ‘exact sciences,’ where some questions at least have exact and verifiable answers. At the opposite end of the spectrum of exactness is sociology, which tries hard to gain respect by using statistics, but to little avail, in my opinion. What we call psychology goes out there, too. Though I personally have found Jung’s work to be very useful, I have no way of conclusively demonstrating that it has done me or anyone else any good. Whether there is even such a thing as effective psychotherapy is questionable in a way that calculus is not.

Concerning Mr Salas’ other question, about attributes that belong to the psyche itself, as distinguished from those that belong to a body-psyche composite animal, such as emotions. I haven’t read far into the text—our initial reading assignment was just the part before the review of previous thinkers, right?—so someone else with a fresh memory of the whole treatise may come straighten me out, that would be welcome—but I’m thinking of dividing our universe of experience into two realms, an inner realm and an outer realm. The outer realm is experienced through the bodily senses. The inner realm is experienced independently of these senses. Pure math and dreams belong to the inner realm. We experience emotions in our bodies as well as in our psyches, and so we consider them to be attributes of the animal in virtue of its possessing a psyche. ‘Recognizes a valid syllogism’ is an attribute of a psyche itself.

16 (Mr Pierce). Mr Porter suggests a distinction, which might be the distinction between emotion and reason. Emotion would require a physical body; reason would not. Perhaps Aristotle alludes to the distinction near the top of 403a, when he mentions thinking. Whether this requires a body remains a question for him.

Mr Salas is apparently worried about asking or at least trying to answer big questions too soon. Certainly we haven’t got much text to go on yet. Mr Salas asks about the first paragraph of the text. A bit later, at the end of 402a, Aristotle asks for the ‘summus genus’ of the soul (if I have correctly rendered the singular of translator Smith’s ‘summa genera’). Aristotle is alluding to the Categories, I take it. I should think that one would want to say what kind of ‘category’ or predicate soul was, before one said much more.

One of the categories, as I recall, is ο�σ�α, translated (perhaps misleadingly) as ‘substance’ or ‘essence’. As Aristotle goes on, he seems to treat soul as an ο�σ�α with attributes or ‘affections’. Does this mean he has answered his categorical question?

At the end of the section, Aristotle says:

the affections of soul are inseparable from the material substratum of animal life, to which we have seen that such affections, e.g. passion and fear, attach, and have not the same mode of being as a line or a plane.

So being fearful is not the same kind of thing as being straight. Now, in the previous paragraph (on 403a), Aristotle says that, if the soul is not capable of separate existence, then it will be like what is straight. Can we conclude then that, for Aristotle, the soul is capable of separate existence?

I am finding it frustrating not to have the Greek. It turns out that Blackwell’s in the UK has the Ross edition of 1956, so I have ordered it. What Greek phrase, for example, does Smith translate as ‘enmattered formulable essences’?

17 (L. Parson). I’m not really following this thread, since I have yet to [get] my Ross out of the box it lives in, but couldn’t resist the opportunity for a bit of pedantry: the phrase you are looking for is ‘summum genus’—‘genus’ is a neuter noun.

18 (P. Lewis). We seem to have touched on this, but allow me reiterate the question in my own humble, simple way… Why does Aristotle use ‘straightness’ to suggest that the soul is inseparable from the body in 403a10–15, but says that these afflictions of soul that are inseparable from body have not the same mode of being as a line in 403b18–20.

19 (Mr Thomas). Well, my first answer to Question one of (11) is that Aristotle didn’t actually claim what the question claims he claimed.

One of the difficulties in reading Aristotle is that we don’t quite know what the words we have were intended for. Aristotle did not publish the works that we have, and a popular notion is that what we have as The Works of Aristotle are mostly lecture notes, prepared either by him or by his students. So let’s assume that what we have as De Anima is just such a set of lecture notes. Accordingly, had we been around way back when we might have come to the lecture series that begins with the sentence at 402a. I guess it’s fair to assume that we would have known (by reading whatever 4th Century flier had publicized the lecture) that we were going to hear a lecture about ψυχ� given by the distinguished Professor Aristotle. What would we have been expecting to hear?

Well, apparently one of Aristotle’s published works (which we might have either read or heard about) was a dialogue on the subject of the immortality of a human being’s ψυχ�. I’m not next to any reference materials, so this is all based on hazy memory, but I think that the scattered reports that survive of the dialogue seem to indicate that Aristotle took a Platonic line in his dialogue; the dialogue was designed, perhaps, to comfort people who were grieving over the death of a friend or family member by giving them grounds for believing that their loved one was not gone forever, but merely relocated someplace.

If we weren’t familiar with Aristotle’s published work discussing ψυχ�, we would probably be familiar, at least by hearsay, with some of the discussions in Platonic dialogues. In the Phaedo, the Phaedrus and the Republic, for example, the ψυχα� of men are discussed. The Phaedo contains an argument for the immortality of human ψυχα� and both the Phaedrus and the Republic contain elaborate reincarnation myths. So it is a beguiling speculation that at least some listeners were expecting to hear more of the same.

So, what does Aristotle start with? The first words out of his mouth are ‘We take knowledge to be fine and worthy, but some knowledge more so than other sorts’—fine and good, and not at all surprising or controversial. But then Aristotle explains why we take some areas of learning to be better than others: ‘EITHER in virtue of its accuracy OR as being of better and more wondrous stuff’—at this point, he has still not asserted anything about whether the study of ψυχ� is more highly prized than other studies for any particular reason. But it does seem to me that, at the beginning of a lecture course on ψυχ�, Aristotle is even by now implying that such a study is more prized than many others. So it occurs to me to wonder what one should think at this point, one quarter of the way through the first sentence.

A way to think about it is to posit^{1} that Aristotle begins by asserting the following proposition:

A field of study is more highly prized than others for one of two reasons: either that field allows for knowledge of exceptional precision and accuracy or that field concerns itself with things that are better and more marvelous than other fields.

If we had heard only that proposition, followed by an acknowledgement that the study of ψυχ� was one of those highly prized studies, what would our expectation be (as educated members of Aristotle’s audience) as to the reason for the high status of the study of ψυχ�?

1I say ‘posit’, because I do believe this proposition is implied by Aristotle’s words.

I would expect that most of us would have thought that the most likely reason to value the inquiry into ψυχ� was because of its subject matter being better and more marvelous than most other subject matters. An obvious example of a study highly prized for its precision, on the other hand, would be geometry. Thus, one can easily imagine that the general preconception would be that the study of ψυχ� is highly prized because the ψυχ� is itself a thing of the better sort, and much to be wondered at. (And I think that this would have been true in 4th Century Athens, and is generally true today.)

Accordingly, Aristotle’s next move is designed to surprise us. Rather than choosing one of the two, he answers, in effect, ‘Both!’ Our modern day surprise leads us to remember the accuracy claim, because it’s the one that doesn’t appear to make sense initially. But we may be wrong to be surprised only at the one claim.

Before I say why this appears to me to be the case, let me advert to Mr Salas’ cryptic caveat in (14), sent subsequently to his opening questions. When I read Mr Salas’ caveat, I was struck by its cryptic nature. It seems utterly true to me, but one would have ordinarily supposed that such a caveat would be issued in conjunction with an explanation of what the term in fact means. I toyed myself with trying to come up with some discussion of what the term ψυχ� means. But when I tried to formulate such an explication, I quickly ran into difficulties. So while I feel strongly that the English word ‘soul’ is too encumbered with a whole squadron of religious and other connotations to adequately translate the Greek term, I don’t have a recommendation as to an alternative (which is why I use the wordlet ψυχ�).

The problem can be made more concrete by fastening onto the second set of Aristotle’s alteration given above: is the ψυχ� ‘better and more wondrous’ than most other items furnishing the world? Well, if we think about ψυχ� as the term is used in the Platonic dialogues we’re familiar with, it’s pretty easy to assert a definitive ‘yes’—after all we all appear to have a ψυχ� and it appears to be quite important in making us the human beings we are. It appears to be the deathless part of ourselves, and immortality is traditionally seen as both better and more amazing than mortality. So thinking in this way, it still appears as if the accuracy claim is the weird claim, demanding explanation.

But wait a minute: as we shall find out a little later, human beings are not the only creatures with ψυχα�. All animals and all plants will turn out to have a ψυχ� as well. This seems to me to make the second claim in Aristotle’s alternation every bit as surprising as the accuracy claim. Is it really obvious that the ψυχ� of a rutabaga (whatever that would mean) is ‘better and more amazing’ than most other entities in the world? Indeed, the claim that the ψυχ� of a gnat is ‘better and more wondrous’ than, say, a trireme could easily be imagined placed as withering satire in the mouth of the Socrates of Aristophanes’ Clouds. Yet, it appears that sober, scientific Aristotle implies just such a claim.

As a logical matter, one could imagine the position that the study of ψυχ� is of major importance because some of it falls under the first alternative and the rest falls under the latter. (Thus, one could claim that the relatively simple and uncomplicated ψυχα� of plants are capable of extreme accuracy, whereas the complex and mysterious human ψυχ� is marvelous indeed.) But this position seems unnatural to me, and the manner of Aristotle’s expression of the situation would not naturally lead to such an interpretation. (It would have been easy enough for Aristotle to say ‘one or the other’ rather than ‘both’.)

Thus, I would reformulate Mr Salas’ question to posit that both alternatives seem mysterious. I think we need to keep this mystery in mind while we explore what Aristotle thinks ψυχ� is.

As to the question of what �κρ�βεια means, Hicks cites to a number of other uses of the term in the other writings of Aristotle, and comments that the term often means ‘abstract’ or ‘involving the first things’ for Aristotle. I haven’t had time to check out the passages Hicks cites, and ‘abstraction’ does seem like a plausible gloss in this passage (although it’s a meaning not given in LSJ^{2}). ‘Concerning first things’ does not seem plausible, because it would then make of the first sentence a circularity, since Aristotle claims that on the basis of both alternative criteria the inquiry into ψυχ� is ‘of the first importance’ (as Hett’s translation has it). The Greek has �ν πρ�τοις which (I’m away from any reference sources, so the real classicists should correct me here) I would say literally means ‘among the first’. If �κρ�βεια was taken to mean ‘involving the first things’ then the sentence would be rendered tautological, and somewhat pointless to say.

Well, I suppose that’s enough for now. I’ll try to move on to the next sentence.

20 (M. Billington). Mr Porter wrote [in (15)]:

I, too, wonder what Aristotle could possibly have been thinking when he starts out asserting that the science of the psyche (dare we call it ‘psychology’?) is to be respected because of its accuracy (or ‘exactness’ in my translation).

(I wouldn’t call it psychology.) I think an answer is in your last line: ‘“Recognizes a valid syllogism” is an attribute of a psyche itself.’ A syllogism is exact. Thinking quâ reasoning strives for exactness. So maybe the edifice constructed by speculative reasoning, i.e., all of his life’s work, is what Aristotle is thinking of.

Mr Pierce seems to have had the same thought [in (16)]:

2‘LSJ’ is the standard abbreviation for the standard Greek-English Lexicon: A Greek-English Lexicon, With A Revised Supplement, 1996 (Ninth Revised Edition), compiled by H. G. Liddell and R. Scott, revised and augmented throughout by Henry Stuart Jones, with the assistance of Roderick McKenzie and with the co-operation of many scholars, Supplement edited by P. G. W. Glare and A. A. Thompson, Oxford University Press.

Mr Porter suggests a distinction, which might be the distinction between emotion and reason. Emotion would require a physical body; reason would not. Perhaps Aristotle alludes to the distinction near the top of 403a, when he mentions thinking. Whether this requires a body remains a question for him.

At 413b25 and again at 430a20 Aristotle has arrived at the position that thinking is separable.

‘Can we conclude then that, for Aristotle, the soul is capable of separate existence?’ It sounds like most parts or powers of the ψυχή, the nutritive, appetitive, sensory, and locomotive, are not capable of separate existence. Being fearful needs a body. Only one part, the thinking part, is capable of separate (and immortal and eternal) existence. Thinking has a mode of being like that of straightness.

Mr Salas began by asking: ‘what do we make of Aristotle’s claim, en passant, that there are some attributes peculiar to the soul and some that the animal shares?’

Here’s a first-paragraph guess, which may or may not stand up: that thinking alone is peculiar to the ψυχ� and that the other powers of the ψυχ� are imbued into, and shared by, the living body.

‘Does Aristotle commit himself to anything by claiming that some attributes are peculiar to the soul?’ No commitment, just a foreshadow of thought thinking itself.

21 (Mr Thomas). Mr Pierce writes in (16): ‘A bit later, at the end of 402a, Aristotle asks for the ‘summu[m] genus’ of the soul. Aristotle is alluding to the Categories, I take it.’

One should be careful in treating such apparent cross-references. It does appear that the text is referring to the work we know as the Categories, but it is quite possible that the original text by Aristotle (or perhaps a student of his) did not include such a reference. Copies of these texts were most likely sometimes annotated in the margins, and modern editors sometimes conclude that bits of such annotations were mistakenly inserted into the text itself in one of the many copies made. (Our evidence for the texts consists of manually copied texts many generations removed from the original, and each generation represents an opportunity for more error to creep in.)

The Categories itself presents a number of puzzles about what it was intended to be. As a result, it may be misleading to import purported doctrine from that text into this discussion.

22 (Mr Thomas). On �κρ�βεια: Ross says ‘It is not very clear why Aristotle assigns a high degree of �κρ�βεια to psychology; Philoponus is probably right in thinking that it is because soul is a pure form, not a complex of form and matter.’ 23 (S. Whitehill). Aristotle is saying in his first line that the soul is worthy of study because being able to accurately describe something is a worthwhile activity in itself with the suggestion that there is a method of looking within, i.e. introspection, that is capable of producing real �κρ�βεια knowledge.

24 (Mr Thomas). Introspection? I’ve just read the whole work and don’t see introspection as a method Aristotle uses at all. Maybe I’ve missed it, but most of what Aristotle discusses are faculties common to all animals, and introspection does not yield insight into what sensation might be for a gopher.

25 (Mr Lewis). Is it introspection to assume that there is a soul in the first place? Not to assume that these passions are merely emanations of the body? Aristotle seems to come to this discourse with some assumptions already made.

26 (Mr Thomas). ‘Is it introspection to assume that there is a soul in the first place?’ I assume you didn’t mean this statement literally, as one can assume all sorts of things without the benefit of introspection.

But you seem to be maintaining, for instance, that it is by means of introspection that Aristotle concludes that mosquitos have ψυχα�. I just don’t see what sort of introspection could possibly yield such a result.

‘Not to assume that these passions are merely emanations of the body?’ I think you will find that Aristotle’s theory doesn’t have the Cartesian mind/body duality that this question seems to assume. For Aristotle, the passions did seem to be physiologically based, but we are not far enough along in the text to really consider the question.

‘Aristotle seems to come to this discourse with some assumptions already made.’ Of course he does. On the other hand, Aristotle tries to lay out his own presuppositions, not always with complete success, of course. But I haven’t even come to the second sentence.

The real danger, as Mr Salas pointed out in his caveat, is that we will bring our own assumptions about ‘soul’ into this discussion of ψυχή.

27 (Mr Lewis). ‘“Is it introspection to assume that there is a soul in the first place?”’ Alright, I realize this is silly but, Aristotle didn’t say that emotive qualities were caused by the gods or the wind or any such thing. He lays it on the soul. I don’t care who wrote this, student or teacher, but in saying that this particular study is of a higher type of knowledge suggests that he has thought about the potential for other causes and that he has ruled them out, or to be lesser, this takes a certain degree of inward looking that I define as introspection.

28 (Mr Thomas). Mr Lewis, can a dog be angry? How can you tell? Must it be by introspection that you deduce a dog is angry? Or perhaps by observation?

More to the point, can you tell us where in the text you are gleaning this?

29 (Mr Lewis). I was just looking at the first lines[.] 30 (Mr Thomas). Here is your original message: ‘Is it introspection to assume that there is a soul in the first place? Not to assume that these passions are merely emanations of the body?’ You seem to be asserting several things here.

1. That Aristotle assumes that there is ‘a soul’. Since he’s starting some sort of lecture course, probably, called ‘Concerning ψυχή’, I’ll readily grant that he is assuming that the word is meaningful. Of course, at 402a23 he says that the first thing to do is to figure out what sort of thing ψυχή is, and he mentions possible kinds of things that it might be that don’t actually include what we would think of as thinghood, such as the possibility that it is a how or a how much or another sort of insubstantial thing. He mentions the possibility that it might be a sort of possibility rather than an actuality (although I think pretty much each English word I’ve used in this sentence is seriously misleading). So in some sense Aristotle does assume the existence of ψυχή, but the mode of existence is not yet specified. I wonder if you think that Aristotle is assuming the existence of ψυχή in some determinate manner? I don’t think he is here. (Of course, later on he will offer his definition, which he thinks does specify the mode of existence. But that’s not in the beginning.)

2. Second, you seem to be asserting that Aristotle can only come up with this assumption about the existence of ψυχή by means of introspection. This I don’t follow at all. To repeat yet again, Aristotle is not talking only about human ψυχα�. He purports to investigate ψυχή in an unrestricted sense. Please do tell me how one can determine by introspection the existence of something that insects have. How does one go about introspecting about mosquitos? Or dogs? Or even other human beings?

3. Where in the first lines of the book are ‘the passions’ mentioned? The word π�θη is used in 402a9, to be sure, but this word does not mean ‘passions’ in the sense we think of them in English. I think we will discover that one of the ‘passions of the soul’ (π�θη τη̃ς ψυχη̃ς) is, for example, sight. If by ‘passions’ you mean to include, for example, seeing a blue ball, then you have to explain what introspection has to teach us about that. If you don’t mean to include sensation as among the ‘passions’, then you are misunderstanding Aristotle. (Or reading a bad translation—a better translation of π�θη is ‘affection’—not in the sense of the emotion but in the sense of being affected by something.)

I fail to see introspection anywhere in the treatise, but I utterly fail to see how it appears in the first paragraph. I wonder if you’re not thinking that Aristotle is engaging in a discussion that would be recognizable to the participants in the Phaedo as the same topic. I don’t think he is. In particular, I don’t think that the question of the immortality of the ‘soul’ is a question that Aristotle even considers for a nanosecond in De Anima.

31 (Mr Lewis). [Concerning (28)]: Another way of saying my ‘silly little point’ is to ask why you assume a dog gets angry in the first place? That, to me, would require a degree of anthropomorphizing that borders on intuition. That isn’t to say that a dog does feel angry, but that my positing it seems to require some ‘6th sense’[.] Aristotle argues that the senses must be separate, for, if it were not like this our perception of the common qualities would always be incidental, i.e. as is the perception of Cleon’s son, where we perceive him not as Cleon’s son but as white, and the white thing which we really perceive happens to be Cleon’s son. But then, what sorts out the white thing from the son thing, and what places this at an order of knowledge that is higher that the perception of white or son?

32 (Mr Lewis). I guess I might have been [‘thinking that Aristotle is engaging in a discussion’ on the topic of the Phaedo (30)], thank you Mr Thomas for taking so much time to indulge me on a point, which, in all honesty, I hadn’t even given a thought until Mr Whitehill mentioned it, and am not sure was worth the energy I made you put forth. Your efforts are quite convincing however and I will gladly put my speculations aside for more ‘worthy’ endeavors to knowledge.

33 (Mr Porter). It would be useful, wouldn’t it, if we all read the whole thing through, at least a quick, superficial read for awareness of content, as preparation for our intended slow reading?

I was going to say, maybe we originally arrived at the concept of a psyche through introspection, and then applied the concept to animals through observation of them. But, I have browsed ahead a little bit, and I want to ask, is it roughly correct to say that the ψυχ� Aristotle and his fellow Greeks discussed was supposed to be the thing that imparts life to otherwise inanimate matter? This would be a concept that could arise out of observation of the world around us, independently of what we observe ‘within ourselves.’

34 (Mr Thomas). Concerning (31): I don’t think that I assume that a dog gets angry. I have seen a number of dogs who act quite angrily. This is called observation, and does not rely upon introspection. Introspection does not even always access my own anger, as I can get angry without realizing it. It has happened on occasion that I required a friend to point out that I was acting from anger to realize that this had happened. But heightened alertness and aggression, together with raised hackles and the other physiological attributes of anger can be observed in dogs with quite relative frequency. So can the anger that arises out of meanness, when you see one dog purposefully making another dog’s life miserable, out of apparent spite. (This is what’s going on in my sister’s household, with a new, retarded adolescent dog being ⟨ahem⟩ hounded by a long established, and much more clever, terrier.)

You are the one making the apparent assumption that anger is exclusively an internal feeling. Of course, I can’t know exactly how a dog feels as he is snarling, with hackles raised and an elevated heart rate. But I can’t know how you exactly feel when you exhibit the signs of anger. So I don’t think that I need introspection to recognize anger either in you or in Fido. And I don’t think I need to ascribe to every sentient being the ‘inner experience’ I access through introspection in order to recognize their anger, for example.

Does a mosquito get angry? I don’t know. Perhaps if I had more experience with observing them I could form an opinion. We do say that bees get angry on occasion, but again, I have too little experience with observing them to make a judgment.

I fully recognize that there are all sorts of puzzles about how one can know things. And you seem somewhat enamored of the Cartesian puzzles. But you also seem to insist that Aristotle is by some sort of necessity a Cartesian thinker, worried about solipsism and trying to bridge the chasm between his personal introspection and the mysterious external world. If you want to insist on this, fine—I’ll just shut up. But you won’t be understanding Aristotle, in my opinion.

35 (Mr Thomas). I think that Mr Porter’s formulation in (33), of ψυχ�, is indeed ‘roughly correct’, at least as far as Aristotle is concerned. Plato’s discussions of ψυχή don’t seem to conform to this rough correctness, though. But then Plato is not trying to give a scientific account of the world, whereas Aristotle probably is. I forget in which dialogue Plato propounds the notion that the ψυχή is the principle of self-motion, a feature of ψυχή that Aristotle also discusses, and a feature that is not peculiar to the human ψυχή. I don’t know enough about how the word was used in other contexts to be very helpful about what Aristotle’s ‘fellow Greeks’ may have understood about the term.

36 (Mr Lewis). [Mr Thomas writes:] ‘I’ll just shut up. But you won’t be understanding Aristotle, in my opinion.’ I am simply trying to understand the agency with which one arrives at the conclusion that certain types of knowledge are ‘superior.’ I do not believe in such a thing, nor do I believe that one can know a dog gets ‘angry’ without anthropomorphizing. If that prevents me from understanding Aristotle, alas, I will read on.

37 (Mr Thomas). ‘I do not believe in such a thing… ’ Do you really mean this? Do you think that knowledge of quantum physics and knowledge of the names of the Spice Girls are equivalent in value? Really?

‘… nor do I believe that one can know a dog gets ‘angry’ without anthropomorphizing.’ I expect that you mean something different by anger than I do. Or maybe you just don’t know dogs.

38 (Mr Lewis). That is true. Most dogs that I know have a terrible time with English.

‘Do you really mean this?’ Yes. On a lighter note, if a crazed fan were holding a gun to my head and were going to kill me if I didn’t profess my love of [C]innamon [S]pice, then I would hold that particular piece of knowledge in the highest regard.

39 (Mr Whitehill). Mr Lewis wrote [in (36)]: ‘I am simply trying to understand the agency with which one arrives at the conclusion that certain types of knowledge are “superior.”’ But Aristotle is not denoting a type of knowledge but an object of knowledge.

The study of the soul is at the front rank of possible enquiries because either it has greater exactitude and a worthwhile object or is ‘more honorable and precious’ than another. Whatever one brings to the De Anima in terms of belief about the soul, its study fits either of these categories and so we should put it among the first rank of the things into which we enquire.

I mentioned ‘introspection’ because it is puzzling how Aristotle will complete his errand by denotation and description alone. And there is the tantalizing question of ‘exactitude’ which opens the door for a discussion of methodology in conjunction with this study. The first question Aristotle asks is how do we study something like ‘the soul’, something which we talk about frequently, a concept that Socrates and Plato denoted in a very thinglike manner. One method of enquiry may not be sufficient to cover all questions.

40 (Mr Lewis). ‘But Aristotle is not denoting a type of knowledge but an object of knowledge.’ That helps considerably! Thank you.

41 (Mr Thomas). One more comment on the first sentence: Aristotle does not literally assert that the study of ψυχή is among the first in importance among sciences. He actually says that it would be ‘reasonable’ (ε�λ�γος) to place it among the first (or ‘highest’) sciences. This may seem like a legalistic quibble, but I think it does emphasize that for Aristotle, these beginnings are still somewhat provisional. His statements about this study being more accurate and better and more awesome are not to be seen as utterly assertoric, but as promissory notes to be redeemed by the end of the course.

42 (Mr Billington). Mr Pierce wrote [(16)]: ‘I am finding it frustrating not to have the Greek.’ Here it is in Greek, with a commentary in Greek:

http://www.mikrosapoplous.gr/aristotle/psyxhs/contents.html On my

browser, I had to click to agree to install the Greek fonts.

43 (Mr Thomas). Concerning the Second question of (11): First, Mr Salas should have put quotation marks around the phrase ‘⟨essential⟩ attributes’, for this, including the angled braces around the word ‘essential’ is what Hett gives in his translation.

Such angled braces are used in classical texts to indicate that the editor believes a word is missing and must be supplied even though there is no textual evidence for the reading. Since Hett’s translation is in the Loeb volume, with a facing Greek text, I took Hett’s use of angled brackets in his translation to indicate that he believes the word should be there (presumably for consistency with Aristotelian doctrine given elsewhere), even though there is no textual reason to include it.

So, Hett at any rate will answer Mr Salas’ parenthetical question: no.

44 (Mr Thomas). Again concerning the Second question of (11): First, Aristotle does not in fact claim that there are attributes peculiar to the soul. He uses the word δοκει̃: ‘it seems’. Aristotle could, in a later stage of the investigation, conclude that there weren’t any such private attributes after all.

Second, in response to this question people have lept straight into the vexed question of the soul’s separation, and implied immortality. Mr Salas urged caution before tackling these questions, a counsel I approve. Nevertheless, an apparently analogous structure occurred to me that might be helpful in sorting out what Aristotle might mean by �δια attributes:

Suppose that we are studying kings. What could it mean to say that some of the attributes of kings are peculiar to a king, while others seem to belong to kingdoms as well, by virtue of the king.

We’ll suppose that the kings we are discussing are absolute monarchs. One of the attributes of such a king is the power to issue binding commands. This is not an attribute of a kingdom, which is a kingdom precisely because it and its parts are subject to the commands of its sovereign. Accordingly, the power of command is an attribute private to the king.

On the other hand, if the king is quick to anger and quick to go to war, this attribute is an attribute both of the king and of his kingdom, by virtue of the king. (A king slow to anger and chary of war will accordingly cause his kingdom to have those opposite characters.)

Note that this says nothing about the independence or ‘separate existence’ of the ruler and his domain. Indeed, a king is only a king so long as he has a kingdom, as Lear learned (to his regret).

I hope that this little story illustrates how it would be possible to have a situation where an entity (here, a king quâ king) dependant upon a larger structure for existence can have an attribute not shared by the larger structure.

By the way, this thought was inspired by the word �ρχ� in the second sentence, which I still hope to get around to discussing. (As a result, this post should perhaps have been postponed until after discussing that sentence.)

I conclude that, both because Aristotle uses the verb ‘seems’ and because the relation he mentions can apply to a variety of states of affairs, Aristotle is not committing himself to any particular consequences when he speaks of �δια attributes.

45 (Mr Salas). I take Mr Thomas’s (henceforth, Mr Th., if it’s OK; I toyed with dubbing you ‘Mr T’ but decided against it) point, that we don’t really know much about what exactly has come down to us in the name of Aristotle. I also subscribe to the lecture-notes hypothesis but can’t say that I’m wedded to it.

The reason that I asked about �κρ�βεια in the first couple of lines is that I wonder about the assumption that this kind of knowledge will be both more wonderous and more �κριβ�[ς] than other kinds of knowledge. Since we have not yet read a definition of the soul, claims for the kinds of knowledge one can have about it are mysterious to me.

Mostly, and in agreement with Mr Th., I think we should keep �κρ�βεια in our minds while we begin to read Aristotle’s definitions of ψυχή. Regarding my caveat, I mentioned it because I think we should be mindful of our translation of ψυχή in order to avoid importing any baggage we may have regarding ‘souls’ into this unusual word. Originally, ψυχή meant ‘breath of life.’ As far as I know, in Greek it can vascillate between mind, heart, and animating principle. In 402a, it seems that we get a definition of the lattermost sort. Ψυχή, in 402a, is the �ρχ� τω̃ν ζ�ων (starting point/principle of animals). As Mr Th. points out, we will shortly read about the various instantiations of the ψυχή. Plants may have them. In general, any living thing may have a ψυχή. Given this rough sketch of ψυχή so far, I have been translating it for myself as ‘animating force’ or ‘animating principle.’ I’ll toss that out as a possible answer to what ψυχή, as a term, may safely mean.

As sheer speculation, I’d say that one thing Aristotle may think is particularly wonderful about ψυχα� is that they seem to be connected with motion. I’m not jumping ahead here but just observing that ψυχα� are associated with animals or living things and motive agency seems peculiar to those living things. At any rate, that’s one of the reasons I think that living things are so amazing.

Mr Th., I haven’t yet read through the entire thread but if you haven’t talked more about �κρ�βεια as abstract, could you? It’s not something that would have come out of the Greek to me but I’d be interested in what you’ve found.

46 (Mr Salas). Mr Lewis, [concerning (25):] I think that we should keep in mind that ψυχή has, as of yet, only been defined as some sort of principle for animal life. Doubtless, Aristotle has come to his treatise with baggage but I don’t think that observing that some objects possess animate qualities and that we can assume that there is a cause for that animation is unfair or unempirical.

[Concerning (27):] We haven’t gotten that far yet. Before we get to π�θη, I think we should deal with book 1. However, it is empirically observable that we have π�θη since we observe each other undergoing things all the time. I don’t think introspection is necessary here.

I think that Ross’ suggestion, quoted by Mr Thomas earlier in the thread is quite reasonable. Insofar as we’ll find that the soul is not a matter/form composite, knowledge of it will be �κριβ�στερον (keener/sharper) than things that are composites for him. Why it is most keen and most wonderous is something about which I’m still wondering.

[Concerning (31):] Whoa there!!

Sensation isn’t for another book. Hold your horses Mr Lewis! We’ll talk about separation a lot, I’m sure, later. One thing that may tide you over until then, though, is that χωρισμ�ς, the adjective translated as separable may suggest logical rather than ontological separation.

Mr Porter, [concerning (33):] I think that’s what we talked about doing. If we hadn’t, I agree with you. We should read the whole but then discuss each part slooooooooooowly. I think we should also try not to jump to parts of the treatise way off in the distance.

‘[I]s it roughly correct to say that the ψυχή Aristotle and his fellow Greeks discussed was supposed to be the thing that imparts life to otherwise inanimate matter?’ I think so.

Mr Lewis, [concerning (36):] Aristotle will later trot out a definition of anger that may satisfy your concern. The definition will have two parts. Let’s call one dialectical or formal and the other material. The material part of ‘anger’ would be blood heating up around the heart. So, this aspect of anger would be defined by physical phenomena. The former aspect of anger, the dialectical one, is (from memory so forgive me if I don’t get it quite right) the response to a perceived slight. So, if I strike a dog and it snarls at me while its blood is boiling Aristotle would claim it [was] angry.

I don’t mean to give you a brief and not-too-engaging response but I know we’ll come to these issues later and don’t want us to lose our way as we move there.

[Mr Thomas, concerning (41):] Good point.

[Concerning (43):] Sorry about being misleading. Since I work with the angled braces all time I didn’t think about it. Thanks for the clarification Mr Th.

That having been said, what do you all think about �σα συμβ�βηκε?

47 (Mr Thomas). Mr Salas writes in (45): ‘I take Mr Thomas’s (henceforth, Mr Th., if it’s OK; I toyed with dubbing you “Mr T” but decided against it).’

Is that pronounced ‘theta’?

48 (J. Tourtelott). ‘“I toyed with dubbing you ‘Mr T’ but decided against it).”’ I pity the fool who do that.

—James B. A. Barracus Tourtelott (check the gold chains)

49 (Mr Thomas). Mr Salas writes in (45):

Originally, ψυχή meant ‘breath of life.’ As far as I know, in Greek it can vascillate between mind, heart, and animating principle.

I don’t think this is accurate. LSJ s.v. ψυχή states ‘Hom. usage gives little support to the derivation from ψ�χω “blow, breathe”’. Readers interested in a capsule version of the standard account of ψυχή can look at the short aticle in the Oxford Classical Dictionary (3rd ed.). Because of the reading reported below, I’m not sure the short article is itself entirely accurate.

I myself am engaged in a diversion from reading the text of De Anima because I became interested in understanding what baggage Aristotle’s contemporaries would have brought to the lecture hall. Accordingly I am reading Toward the Soul: An Inquiry into the Meaning of ψυχή before Plato by David B. Claus (Yale University Press, 1981). When I’m finished with it I will try to give a brief report.

Mr Salas asks:

Mr Th., I haven’t yet read through the entire thread but if you haven’t talked more about akribeia as abstract, could you? It’s not something that would have come out of the Greek to me but I’d be interested in what you’ve found.

I haven’t found anything other than Hicks’ words. I’m wondering, though, if by ‘abstract’ Hicks means more or less what Ross said, that is, form alone has more �κρ�βεια than form with matter.

(Note re: baggage—the ‘form/matter’ distinction in Aristotle is in no way the same as the ‘mind/body’ duality we moderns are stuck with.)

Mr Salas writes: ‘That having been said, what do you all think about �σα συμβ�βηκε?’ It’s been a while since I looked συμβα�νω up. I had forgotten that it’s first meaning was ‘stand with the feet together’.

More to our point, it means to meet, to come together, to make an agreement, to happen, to occur, to result, and a zillion other more or less related meanings. LSJ has a separate section for philosophical meanings, and cites Aristotle for the following two opposing notions:

1. ‘a contingent attribute or “accident” (in the modern sense)’.

2. ‘an attribute necessarily resulting from the notion of a thing, but not entering into the definition thereof’—Aristotle’s example is the fact that the angles of a triangle equal two right angles.

While most would say that this is one of Aristotle’s technical terms, I’m inclined to say that in opposition to φ�σις and �υσ�α, �σα συμβ�βηκε περ� α�τ�ν means ‘whatever else happens to come with it’. The term ‘attributes’ probably says this as well as anything, except to the extent that it appears to be definite about what sort of being an attribute has; I think Aristotle uses it here primarily as a place-holder.

50 (Mr Salas). Συμβα�νω has an impersonal use, which is the use I take this instance of the verb to be. In conjunction with �σα, I think the phrase �σα συμβ�βηκε amounts to ‘such things as happen to be.’ The περ� α�τ�ν amounts to ‘around’ or ‘near’ it. A smoother translation, with περ� α�τ�ν, would be ‘attributes.’ What I wonder at is whether we’re talking about essential or incidental attributes here.

The lack of specificity makes me think that we’re either talking about incidental attributes or Aristotle is not making a sharp distinction at the moment.

51 (Mr Thomas). Mr Salas would apparently translate �σα συμβ�βηκε περ� α�τ�ν as ‘such things as happen to be around it’ (‘it’ = ψυχ�). I like the notion of ‘being around’ here (it’s the major sense of περ�, at least with the accusative).

Since this phrase announces a secondary search after the ones for φ�σις and ο�σ�α, the notion is that these ‘attributes’ are at the periphery, whereas φ�σις (derived from the verb for blooming) and ο�σ�α are almost certainly at the center.

This suggests that the answer to Mr Salas’ question: ‘What I wonder at is whether we’re talking about essential or incidental attributes here’ cannot be Hett’s answer. That is, these attributes are either accidental (or incidental) or ambiguously either accidental or essential. One concludes that Hett’s ‘⟨essential⟩’ is simply wrong.

52 (Mr Porter). I suppose this question is mainly for Mr Salas, who proposed De Anima, but any good thoughts from others are also welcome.

Why are we reading this, not another of Aristotle’s works?

53 (Mr Thomas). I can’t tell you why Mr Salas suggested this reading, but I can tell you why I joined in.

As unlikely as it seems, Aristotle was the Greek philosopher known to the scholastic era, not Plato. (Only the Timaeus was translated into Latin, so all most scholastic scholars knew only what Aristotle claimed were Platonic doctrines, not his actual work.)

This has led to certain oddities, since Aristotle is a relatively secular personality, whose universe and ‘god’ are both uncreated and eternal on both sides of the time line. Plato’s apparent mystical turn is almost gone in Aristotle. But nevertheless, Aristotle’s works formed the basis of scholastic learning.

As a result, even the best translations tend to preserve the technical Latin vocabulary used by the schoolmen. As Michael Frede told me once, we won’t really understand Aristotle until his works are re-edited from scratch, and the scholastic encrustations are eliminated.

I had that conversation with Frede a quarter century ago, right after taking a seminar with him in which he came up with a brilliant interpretation of the Categories in which that work was an early ontological treatise, subsequently bent out of shape to appear as a logical work. (I’m told Frede no longer interprets the work thus.)

I concluded that to read Aristotle well would be almost impossible, and subequent life left me without rereading any for that quarter of a century.

In part, this was because I was engaged (fitfully) with trying to figure out Plato, a difficult enough task on its own.

De Anima is one of Aristotle’s most difficult works, and it involves the very difficult notion of �ντελ�χεια. It also has been claimed by the schoolmen for a proof of the soul’s immortality (a proof far from Aristotle’s purpose, in my opinion), and more recently by phenomenologists as a precursor. In other words, it is a text which has for at least a millenium been subjected to special pleading in its interpretation.

I think it would be a worthwhile project to try to figure out what it says.

I note in conclusion, that Mr Salas proposed this as a text for a ‘close reading’— by which both he and I, I think, mean (as he put it) sloooooww. It is possible to loose sight of the forest when examining the trees, to be sure. But the forest is in fact made up of trees, and thus to figure out the forest one must examine the trees. This is why so many electrons have been spilled on the question of what the phrase translated as ‘attributes’ means.

54 (Mr Lewis). Where should we be in the reading at this point in the discussion?

55 (Mr Thomas). Well, I’m not to the second sentence yet.

The portion we were assigned to discuss was the first two chapters of the first book. (My text is in checked luggage, but I think this means through 405b30, although that could be a wayward memory.)

I think that the opening part needs very slow consideration. I would expect the pace would pick up a bit during Aristotle’s discussion of his predecessors’ views, and slow down again when we reach book 2.

56 (Mr Whitehill). Mr Porter wrote: ‘Why are we reading this, not another of Aristotle’s works?’

I don’t know why Mr Salas picked it but I remember De Anima is where Aristotle’s ontology from the Metaphysics is applied, making De Anima the reading of recourse to aid in one’s understanding of the vital concepts in the Metaphysics.

57 (G. Squires). Mr Thomas: Thanks for your response [(53)] to Mr Porter’s question, ‘Why De Anima?’ It is helpful to know that we’re going to tackle the most difficult of Aristotle’s works! Would you mind elaborating a bit on this word [�ντελ�χεια]?

58 (Mr Salas). I don’t know that De Anima is the most difficult of all Aristotle’s works. I’ll explain my own interests in it in another email, addressing Mr Porter’s question.

As for �ντελ�χεια, the term is usually translated ‘actuality.’ Often ‘actuality’ will also be given as a translation for τ�λος. I’m interested in seeing if there is any consistent use of each in the De Anima. As far as I know the distinction is a tough one, if it’s there. ‘Actuality’ is normally paired with ‘potentiality’ δ�ναμις in Aristotle. We will doubtless discuss Aristotle’s use of these terms as we move through the text. An ad hoc example may just be the potential/actual energy distinction many of us were taught in middle school. I don’t know if this example works as well as I’d like but that’s the rough idea.

59 (Mr Salas). Dear Mr Porter, I prefer De Anima in part because it’s very short. I would have also been interested in the Metaphysics but that’s gargantuan. I also would have liked both Analytics but I don’t know that I’m able to helpfully guide us through them. Actually, I know I can’t. At the time, someone, who incidentally is not in this thread at all, complained that we don’t have sustained discussions on texts. I suggested some possibilities and people voted up De Anima.

Less practically, I’m interested in so called Aristotelian psychology. We’ll have to pull in various minor works and passages from other Aristotelian works to explain a lot of what’s going on in De Anima and I thought that would be neat. As an addendum to Mr Th.’s great explantion, I would add that the Aristotelian notion of perception is central to his epistemological commitments since he can be very empirically oriented when he approaches understanding. De Anima, along with some minor works[,] has the most extended account of perception in the corpus.

60 (Mr Salas). Dear Mr Lewis, I think we’re still on the first few sentences. The way I envisioned this, although I don’t see myself as the exclusive architect of our plan, was to discuss until such time as a couple of people voiced a desire to move along and everyone’s concerns were addressed.

Since I have a tough time devoting tons of attention to the list for swathes of a normal week, I will probably post most in fits and starts.

61 (Mr Thomas). Ms. Squires writes:

Thanks for your response to Mr Porter’s question, ‘Why De Anima?’ It is helpful to know that we’re going to tackle the most difficult of Aristotle’s works!

It’s probably not the most difficult. Some chapters of the Metaphysics probably are. I only claimed it was among the most difficult.

Would you mind elaborating a bit on [�ντελ�χεια]?

I don’t think we are in the right place to do so. �Εντελ�χεια is regarded as one of the crucial Aristotelian terms, and it is apparently a word that Aristotle made up. When we get there we’ll have to spend a lot of time puzzling out what it means.

I’m willing now only to assert that it can’t be ‘actuality’ (the traditional translation), since Aristotle would not have had to make up a word for such a concept.

62 (Mr Salas). Dear Mr Thomas, [Concerning (55):] I would have thought so too but I had a seminar on De Anima where a lot of interesting progress was made by examining what Aristotle retains from his predecessors and tosses out.

We may not want to take even that dry section for granted!

63 (E. B.). Mr Salas suggests:

An ad hoc example may just be the potential/actual energy distinction many of us were taught in middle school. I don’t know if this example works as well as I’d like but that’s the rough idea.

Perhaps, but the distinction is between potential and kinetic. Both are ‘actual’ in a manner of speaking, though perhaps not in the sense in which Aristotle means ‘actual’.

64 (P. Goldsmith). In my innocence, I always thought of �ντελ�χεια as the property a thing may have of containing within itself the thing’s own purposes. Now that it’s rolling around in the brain, I see this is not a simple notion at all.

Back to lurking on this thread. I haven’t read De Anima since 1973 so I’m in grave peril of making a fool of myself.

65 (Mr Salas). Ms B., I thought I had it down once and have since discovered that the distinction in Aristotle can be difficult for me to grasp. At any rate, in Aristotle a thing that exists potentially still has a form of existence.

66 (Mr Thomas). Indeed, at 412a6-11, Aristotle reveals that matter (a) exists,

(b) is not an individual thing, and (c) is potential. And then we find out that �ντελ�χεια has two senses.

But let’s wait until we get there to figure out what the hell he could mean by that.

67 (Ms Squires). Ms B. wrote: ‘… the distinction is between potential and kinetic’.

Mr Salas wrote:

I thought I had it down once and have since discovered that the distinction in Aristotle can be difficult for me to grasp. At any rate, in Aristotle a thing that exists potentially still has a form of existence.

In bringing this concept discussion back to question one, ‘What does Aristotle mean… knowledge of the soul is highest in virtue of its accuracy… ’

Is potential existence accurate, exact, precise? Or does it become so only when it is actual, realized, kinetic? We use the expression that ‘thoughts form and take shape’ in our minds. We write books that are potential books until they are complete and printed. So, how can the study of the soul be accurate?

Is the suggestion that the soul is potential rather than actual? Or is Aristotle setting out to prove that the soul is actual?

Aristotle proposes that knowledge of any kind is valuable. Some kinds of knowledge are more valuable than others. What makes the knowledge of greater value can be one or both of two things,

• accuracy/exactness or

• dignity/wonderfulness of the object.

On both accounts:

(i) all knowledge is good

(ii) some kinds of knowledge are better than others

I’m still thinking that it is not the exactness and accuracy of the study of the soul which makes it a better subject for study, but the wonderfulness and dignity of the subject itself.

I think I’m waiting to be shown that the study of the soul can be an exact science.—But then there’s the 21 grams measurement… that’s pretty exact, isn’t it?

68 (Ms B.). Mr Salas writes:

I thought I had it down once and have since discovered that the distinction in Aristotle can be difficult for me to grasp. At any rate, in Aristotle a thing that exists potentially still has a form of existence.

I have enough trouble with energy, so I won’t venture too far into Aristotle. It might be interesting if energy were a good analogy for what he is talking about, but I would be wary of importing something very specific, also called ’potential’, into a discussion of Aristotle’s terms.

In a side note, I am currently having conversation practice with someone who is just learning English. The other day she wanted to know what ‘actually’ meant, because she hears it used so often and couldn’t figure out what it means. It was extremely difficult to explain, once I started realizing in how many different senses we use it.

69 (Mr Tourtelott). ‘Actually,’ in conversational English, almost always does not mean ‘actually,’ and ‘literally’ virtually never means ‘literally.’

70 (Mr Salas). Ms B., not to get into it too deeply but I think, in fact, that kinetic and potential distinctions derive from Aristotle. I wonder if that’s an old classicist’s tale, though.

71 (Ms B.). Mr Salas, I suspect you are right, at least partially. ‘Vis viva’ probably comes into it somewhere.

72 (L. Travis). I have been trying to answer this question for myself; ‘Why De Anima?’

I loved Aristotle in SJC expressly because the readings did not involve poetry and seemed to fit the way I examine problems. Mathematical steps rather than a more holistic or artsy method.

To me, De Anima was/is important because I have wondered since this reading ‘Why postulate a soul?’ when attempting to understand the ‘way things are in the world’. The reading seems to me to be related to the other more clinical examination of things ‘animal’. De Anima discusses something about which [I] remained facinated, including discussions in Pascal, and later Germans concerning potential and actualization.

Of course, now that we are in the reading and I am reminded of the difficulty of the text… I wonder if I am up to the task.

73 (Mr Porter). The reality of the line by line analysis occurring here is starting to sink in for me, and I’m pleased to think I’ll have a chance to actually understand a lot of the treatise. Are we ready to proceed to the next passage?

‘To attain any assured knowledge about the soul is one of the most difficult things in the world… ’

74 (Mr Salas). Mr. Porter urges us onwards and upwards!

Before we move on, though, I have another question about the status of our ‘attributes’ in line 402a9.

I noticed, this evening, something that might inform our translation of �σα συμβ�βηκε περ� αυτ�ν. The following line, 402a9, glosses these attributes as ‘[Attributes], some of which seem to be affections (π�θη) of the ψυχ� exclusive to it, while others seem to belong also to animals on account of that ψυχ�. Hett translates our �σα συμβ�βηκε περ� αυτ�ν as ⟨essential⟩ attributes.

Mr. Th. and I haven’t seen any textual basis for that translation so far. I wonder, though, whether line 402a9 requires that �σα συμβ�βηκε περ� αυτ�ν be essential attributes?

I think that attributes that belong exclusively to a thing will be essential attributes (or defining attributes) of a thing. I suspect that’s where Aristotle is going here but I also have serious doubts about my suspicion. For example, if I happen to be the only creature in existence with this particular color of green in my eyes, should I be defined by that attribute? I don’t think so and I think that Aristotle would call the color of my eyes an incidental attribute. On the other hand, Aristotle might argue that the particular shade of green my eyes possess is not �δια to them, except incidentally. What I mean is that I just happen to have this color of green but am not required to have it. If that’s his tack here, then the ‘π�θη �δια’ would seem to be essential attributes.

The π�θη that belong to the body also but on account of the soul are also interesting. I think it might work in a way analogous to this: one can say my hand is moving as I walk down Waggener hall. One can also say that my hand, although moving, is only doing so on account of my whole body (or to whatever you want to ultimately ascribe my motion). These π�θη, although shared with the animal (as a whole), arise on account of the soul. I’m not clear on what their status, as incidental or essential, is[,] but Hett may well be taking these two clauses to suggest that Aristotle is talking about essential attributes in the preceding line. What do you guys think?

I’m also not sure, though, of what π�θη means in line 402a9 and whether or not that would affect my above example (of the eyes). Clearly, these attributes are some sort of thing that the soul undergoes or experiences but I’m a little unclear about what exactly A. is talking, while attributes that are shares with other things. Perhaps we can move a little by trying to elucidate this issue: what exactly is A. looking for in regards to the soul?

In 402a7-9, we have the soul’s φ�σις (nature), its ο�σ�α (being/definition), and its attributes as objects of investigation. Shall we consider, in addition to the questions above, what a thing’s φ�σις, ο�σ�α, and attributes are? Caveat: we may have to jump about in other Aristotelian texts to nail down his precise usage but it may be worthwhile considering this laundry list forms the goal of A’s investigation as stated.

Mr. Porter, the translation you’re using, ‘To attain any assured knowledge about the soul is one of the most difficult things in the world’, includes ‘assured’ in regards to belief (π�στις). I don’t see that in the Greek, do any of you?

75 (Mr Thomas). Mr. Salas writes: ‘I think that attributes that belong exclusively to a thing will be essential attributes (or defining attributes) of a thing.’

Why? Take the example of an army. The army includes its general, of course, both as a part and as the ruler of the whole. Some attributes of the general will be attributes of the army as a whole (decisiveness, say, or spiritedness). Other attributes of the general may be exclusive to the general – suppose the general has a delicate stomach, and must carefully watch what he eats. This attribute is by no means an essential attribute of a general, even though our particular general has it. Other generals may have robust digestive systems, and be just as much generals as the first. And despite the adage that an army travels on its stomach, armies don’t in fact have stomachs, even though each member of it does. So the attribute of digestive disposition is one that structurally cannot be held by an army.

76 (Mr Salas). Because the digestive example is not exclusive to the general. By �δια I think that A. means necessarily exclusive.

I’m running off to class right now but is the rest of your example aimed at the attributes that exist on account of the soul?

77 (Mr Thomas). This seems to beg the question. If you think �δια means ‘necessarily exclusive’, then the question is answered. But clearly this is not the ordinary meaning of the Greek word. So if Aristotle uses it in a (pun intended) idiosyncratic way, one would have to demonstrate that.

My example was intended to show a way in which one could say that an attribute could be private to a part of a composite entity without being essential to the part. It was not intended to be a perfect analogue to the ψυχ� as a part of an animal. (Although I suspect it will turn out to be a pretty good analogue.)

78 (Mr Thomas). Mr. Salas writes:

Mr. Porter, the translation you’re using, ‘To attain any assured knowledge about the soul is one of the most difficult things in the world’, includes ‘assured’ in regards to belief (π�στις). I don’t see that in the Greek, do any of you?

Hett has ‘sure belief’, which is better than ‘assured knowledge’. Π�στις is much stronger than the English word ‘belief’. English ‘belief’ is generally agnostic about the truth of the matter believed, whereas π�στις seems to imply a greater degree of confidence in the assurance in the matter concerned. LSJ does not give ‘belief’ as a meaning of π�στις, using terms like ‘trust’, ‘faith’, ‘confidence’, ‘guarantee’ and ‘assurance’. Δ�ξα is the term that would be used for ‘mere belief’.

79 (Mr Lewis). Has anyone read the Thomas Aquinas commentary on De Anima? Is it worth a peek?

Third Sentence:

(i) The soul is a substance (ο�σ�α) that has its own particularity or Distinctiveness but as such can it be suited to any body or just to a specific one?

(ii) By connecting the soul’s principles to ‘animal life’ is Aristotle necessitating an empirical approach to derive the generic concepts from the individuals, which came first?

80 (Mr Thomas). Mr Salas in (74) writes two sentences in the same paragraph:

The π�θη that belong to the body also but on account of the soul are also interesting… These π�θη, although shared with the animal (as a whole), arise on account of the soul.

These two statements may illuminate one of the pitfalls that we will be facing. It is natural to us post-Cartesians to think of a soul/body duality. Thus, the two sentences I have quoted above may appear to be making the same distinction. It is important to note, however, that Aristotle does not claim that some of the affections of the ψυχ� ‘belong to the body also’. When he talks about affections not peculiar to the ψυχ�, he talks about how they belong to the animal on account of the soul. Thus, the second of Mr Salas’ sentences quoted above is a correct paraphrase of Aristotle’s words, whereas the first sentence is incorrect.

I expect that Mr Salas wrote the first quoted sentence without thinking through the difference between an animal and its body, because that is a distinction that is not at all alive to us. But it may in fact be a distinction that is alive to Aristotle, and we should be careful not to confuse the two.

81 (Mr Salas). Mr Thomas wisely cautions us against falling into the Cartesian mind/body trap. In the case that I had mentioned, of movement, I do think it’s useful to talk about a distinction between ψυχ� and body (ςω̃μα) but I should have been more careful in my choice of examples. I didn’t intend to separate the body from the soul ontologically or to suggest that Aristotle would be doing that. I think, however, that movement will belong to the body incidentally as well as to the animal as a whole. I don’t think that implies the animal’s constituent parts are ontologically separable, although it doesn’t rule it out either.

Perhaps a better example is life. Life is clearly a π�θος of the whole animal on account of the soul. If we accept that the soul is the life principle, this example may be a paradigm one.

Incidentally, I’ve been thinking about �δια π�θη this past week and think I’ve come to largely agree with Mr. Th.’s correction of me. I don’t think that these π�θη are essential any longer but there is something peculiar, no pun intended, going on with them. First of all, they suggest that Aristotle does have some kind of ψυχ�/body distinction in mind, although the sort of distinction he’s got in mind is wide open. Second of all, they suggest that the whole animal and its soul may not correlate at all points. I think we should keep an eye out for what Aristotle does with these �δια π�θη in the upcoming pages.

Again, what do you all think about setting up φ�σις, ο�σ�α, and �σα συμβ�βηκε up as objects of inquiry? Do we all agree on what these three terms mean? Should we talk about it?

Mr. Lewis, I’m not sure what you mean [in (79)] by the third sentence. Do you mean ‘The soul is a (or the—unclear) principle of animals’?

Regarding your question, Aristotle will give a straight answer on that one later on when he starts trashing his predecessors.

I’d watch out for the principles ‘of’ the soul. The soul is the principle of animals. It is unclear to me whether the soul could have a principle since it seems to be one. The remainder of your question is what confused me. Could you elaborate?

82 (Mr Thomas). What does ψυχ� mean?

In (49) I reported that I was reading a book^{3} in an attempt to understand what Aristotle’s contemporaries might have supposed ψυχ� meant. I’ve finished it, and it has given me a lot to think about. In what follows I will attempt to summarize those of Claus’ points that seem most relevant to De Anima. I should note that, while I expect that Claus is correct (in part because he seems a judicious and careful scholar), I have not done any research independently and rely almost entirely upon Claus in what follows.

Here’s Claus on the outlines of the story he will tell:

In Homer, ψυχ� signifies both the ‘life’ that is lost at death and ‘shade’ or ‘wraith’… After Homer ψυχ� undergoes transformations in meaning that lead eventually to its use by Plato to designate the comprehensive personal ‘soul’—the immortal and divine part of man, the self as a center or microcosm of his whole being, the seat of the rational intelligence and thus of moral choice, that which is not body and which is related to body as master is to slave.^{4}

3Toward the Soul: An Inquiry into the Meaning of ψυχ� before Plato by David B. Claus (Yale University Press, 1981).

Thus Plato’s ψυχ� looks very much like our word ‘soul’. The question is whether Aristotle follows such a usage in De Anima. I don’t think he does.

So, let’s start with Homer. Claus discusses a number of Homeric terms for ‘life-force’, of which ψυχ� is one. Among these are θ�μος, �η̃τορ, κ�ρ, μ�νος and κραδ�η. Claus surveys the use of all these words in Homer, and notes that ψυχ� is generally more restricted than the others (some of which can be used in personifications, or attributed to animals). Claus speculates that the ‘shade’ meanning of ψυχ� keeps it from being used with equivalent freedom. So let’s first look at ‘shade’ (or ‘ghost’) uses of the term in Homer.

In Book 23 of the Iliad, the ψυχ� of the dead Patroclus visits Achilles. It’s clear the visit starts when Achilles is asleep; it’s not clear to me when Achilles wakes up, however. At 23.62–68 the coming of Patroclus’ ψυχ� is narrated. That ψυχ� is described as ‘in all things like his very self, in stature and fair eyes and in voice, and like were the clothes that he wore about his body’.^{5} Patroclus then makes a speech to Achilles (23.69–92), to which Achilles (still sleeping?) responds (23.93–98).

The next bit is worth quoting in full:

So saying he^{6} reached out with his hands, yet clasped him^{7} not; but the spirit like smoke was gone beneath the earth, gibbering faintly. And Achilles sprang up in amazement, and struck his hands together, and spoke a piteous word: ‘Well now! Even in the house of Hades there is something—spirit and phantom—though there is no mind at all; for the whole night long has the spirit of unlucky Patroclus stood over me, weeping and wailing, and charged me concerning each thing, and was marvellously like his very self.’ (23.99–107)

Of course the most extended sequence with ψυχα� (best translated in this context as ‘ghosts’) is Book 11 of the Odyssey, the sequence where Odysseus calls forth the ghosts of the dead.

In order to call them forth, he sacrifices some sheep and catches their blood in a trough. It appears that the shades come in droves as they yearn to drink the blood. A couple of the shades seem able to speak without first drinking the blood, although that’s not completely clear to me. Odysseus unsheathes his sword to keep the shades from the blood until he has had his chat with the ghost of the prophet Teiresias.

4p. 1.

5Quotations from the Loeb Second Edition (1999), translated by A.T. Murray, revised by William F. Wyatt.

6Achilles.

7Patroclus.

Among those Odysseus sees is the ghost of his mother, who died while he was at Troy. Yet she appears not to recognize him. So he asks Teiresias why this is so:^{8}

‘I see here the ghost of my dead mother; she sits in silence near the blood and cannot bring herself to look upon the face of her own son or to speak to him. Tell me, my lord, how she may recognize that I am he?’ So I spoke, and at once he made answer and said: ‘Easy is the word that I shall say and put in your mind. Whoever of those that are dead and gone you shall allow to approach the blood, he will speak truly to you; but whoever you refuse, he will go back again.’

So saying, the ghost of the lord Teiresias went back into the house of Hades, when he had declared his prophecies; but I remained there steadfastly until my mother came up and drank the dark blood. At once then she knew me, and with wailing she spoke to me winged words: (11.141– 154)^{9}

After Odysseus and his mother chat for a while, Odysseus tries to hug her ghost. ‘Three times I sprang toward her, and my will said, “Clasp her,” and three times she flitted from my arms like a shadow or a dream.’ (11.204–209) He then asks whether she’s a phantom (ε�δωλον) sent by Persephone to torment him. She answers:

Ah me, my child, ill-fated above all men, it is not that Persephone, daughter of Zeus, is deceiving you, but this is the appointed way with mortals, when one dies. For the sinews no longer hold the flesh and bones together, but the strong force of blazing fire destroys these, as soon as the spirit (θ�μος) leaves the white bones, and the ghost, like a dream, flutters off and is gone. (11.216–223)

A little later Odysseus chats with the ghost of Achilles, who rather famously says:

Never try to reconcile me to death, glorious Odysseus. I should choose, so I might live on earth, to serve as the hireling of another, some landless man with hardly enough to live on, rather than to be lord over all the dead that have perished. (11.488–491)

8In this section of the Odyssey Odysseus is narrating his past adventures to the Phoenician court, so technically there should be quotation marks within quotation marks to indicate that we are listening to Homer recounting what Odysseus says about what he did and what he heard from others, with those quotations from others receiving in turn their own quotation marks.

For the purposes of these quotations, however, the outermost quotation marks (sometimes two sets of them) will be silently omitted.

9Quotations from the Loeb Second Edition, translated by A.T. Murray, revised by Geor[g]e E. Dimock (1995).

In short, whatever existence the ψυχα� of the dead have, it is an inhuman, uncanny and unpleasant existence. Frequently the dead are given the epithet ‘strengthless’, and they appear not to be functioning or rational without some infusion of strength from elsewhere (the sheep’s blood, for example).

It’s hard to imagine anything further from the Odysseus in the final myth of Plato’s Republic, calmly and rationally choosing his next life.

Of course, every educated Greek would have been quite aware of the Homeric use of ψυχ�.

But what happened to the word in the period from Homer to Aristotle? Claus goes into the history through Plato. He finds that the word comes more and more in popular usage (by which he means usage by folks other than philosophers and doctors) to carry the generalized meaning ‘life-force’. (Most of the other Homeric words for ‘life-force’ go out of the language. The exception is θ�μος, which comes to mean ‘spiritedness’, and which Plato in the Republic makes a part of his tripartite ψυχ�.)

Claus finds, more surprisingly, that even in technical uses (by Pythagoreans and medical writers) the ψυχ� is never regarded, before Plato, as a psychological agent, directing the body independently. This Platonic use Claus opposes to what he calls the ‘psychosomatic’ notion of ψυχ�, in which the ψυχ� is either a physiological part of the body or is the means by which the body (and thus the person) is affected.

Thus, it would appear that Aristotle, in bringing the ‘life-force’ meaning to the fore, is (broadly speaking) talking about the common notion of ψυχ�, rather than the weird notions that Plato came up with.^{10} Thus one can see in Aristotle’s approach in De Anima a rather significant breach with his teacher Plato.

There are a few odd things, though, about the term ψυχ�. According to Claus, ψυχ� is never used in Homer with reference to animals, and its use with reference to animals elsewhere is apparently qwite rare. (Homer often uses the other ‘lifeforce’ words with reference to animals.) Thus, it would appear that Aristotle’s generalization of ψυχ� to non-human realms would have been an innovation. (An alternate explanation would be that few pre-Aristotelian writings concerning natural history have come down to us, and that this use is accordingly not attested in the record but did in fact exist.)

Hope this has been helpful. Now on to sentence 2!

83 (Mr Thomas). Mr Salas writes: ‘Perhaps a better example is life. Life is clearly a π�θος of the whole animal on account of the soul.’ It’s far from clear to me that ‘life’ is a π�θος of an animal. I would have said that an animal without life is in fact not an animal at all. (Or ambiguously so—as in Monty Python, one could say that a dead animal is in fact an ex-animal.)

A π�θος seems to be something that happens to some subject, but when life departs a living thing, the living thing is itself destroyed, not changed as a continuing subject.^{11}

10Remember, though, that Aristotle apparently published a dialogue using a Platonic notion of ψυχ�.

84 (Mr Whitehill). Mr Salas wrote:

I’d watch out for the principles ‘of’ the soul. The soul is the principle of animals. It is unclear to me whether the soul could have a principle since it seems to be one.

This is what the text will explore although I agree that the soul will turn out to be the principle of animal life. Aristotle will be examining a ‘black box’ in terms of its manifestations, and look into the manifestations to tease out its properties. Aristotle is able to successfully carry out a τοδε τι of soul and it is meaningful, but it’s analytic in a similar form to those Euclidean propositions that are regarded as analytic, with the statement ‘let it have been done’ and the demonstration proceeds to examine geometric properties. Aristotle says at the outset ‘there is’. In other words, he posits the existence of soul. If it exists, then it is knowable by its manifestations and distinctions from mere matter. Does a rock produce behavior? No. Then what is the principle by which matter can produce behavior?

We’re no where near the Cartesian identification of the soul as a separable entity. We’re to examine the difference between matter (�λη) and life (ζω̃�ον). Descartes’s cogito is the affirmation that the soul has a kind of ‘life’ of its own, but for Aristotle, the soul may be the essential principle of life. An interesting question to ask as we read is what does a rock lack to be potentially a life?

85 (Mr Salas). Dear Mr Thomas, Although I’m not yet clear on what limits there are on π�θη, I know that Aristotle admits of active ones (on the next page of our reading actually). Since sensation, desire, and anger all qualify as π�θη, I wonder whether the whole animal experiences life as a π�θος.

Thanks for your post on the history of ψυχ�. It was interesting.

What do you and what do you all think of that tripartite question, concerning φ�σις, ο�σ�α, and �σα συμβ�βηκε?

86 (Mr Thomas). Mr Salas writes: ‘Since sensation, desire, and anger all qualify as π�θη, I wonder whether the whole animal experiences life as a π�θος.’

The affections seem to be particular.^{12} My seeing, which I do more or less continuously, is always a seeing of something. I see a red ball, or a yellow taxi, or a baseball game. What my affective seeing is not is generic: I have never seen in general, I’ve only seen particular sights.

11Since π�θη is plural, it does not seem to be grammatically possible for ‘life’ to be a π�θη, but it doesn’t seem to make sense for ‘life’ (as an abstract term) to be one among a number of π�θη of an animal either. (This footnote, and the corrections in the form of the Greek in the post itself, was added after the original post, when I realized that I was missing the distinction in Greek between the singular π�θος and its plural.

12The original post had ‘singular’, which upon reflection is not quite what I had in mind.

It also reads oddly in light of the muddle I made (which I’ve tried to fix in this transcript) of confusing the singular and plural forms of π�θος.

It’s not clear to me that ‘life’ isn’t like seeing in general, and accordingly not like seeing (or desiring, or knowing) at all. Thus my scepticism about whether ‘life’ is itself a π�θος.

87 (Ms Squires). Using this online dictionary

http://www.kypros.org/LearnGreek/mod/resource/view.php?id=475 I am

able to only find ο�σ�α = essence, nature, savour, substance

What is the meaning or the English translation of the other words [in (85)]?

88 (Mr Salas). The usual translations for φ�σις, ο�σια and �σα συμβ�βηκε are nature, essence, and attributes respectively.

89 (Mr Pierce). In the first three sentences of the text, φ�σις appears twice. I think the word has two meanings, just as the English word ‘nature’ may refer either to living things collectively, or to the supposedly unchanging properties of a particular thing. I want to say that this second meaning of ‘nature’ is ‘essence’.

Indeed, Aristotle first says that knowledge of soul appears to contribute greatly to the pursuit of any truth, especially in φ�σις, since soul is a sort of principle of living things.

We seek, says Aristotle, to know the φ�σις and ο�σ�α of soul, and then �σα συμβ�βηκε.

In the latter passage, Smith treats the φ�σις κα� ο�σ�α as a ‘hendiadys’, if that is the term, a one-thing-through-two-words: ‘essential nature’. Fair enough, I suppose; then Aristotle is making not a tripartite, but a bipartite division.

If I understand the Greek correctly, �σα συμβ�βηκε means literally or etymologically ‘as much as has stood together with’.

How does one parse the ε�ι̃θ� that precedes this term at 402a8?

90 (Mr Thomas). This is the adverb ε�ι̃τα, which just means ‘then’ or ‘next’. It has lost its final vowel, and the tau becomes a theta due to the fact that the following word carries a rough breathing. LSJ notes that it is frequently used to express surprise, incongruity, indignation, contempt, sarcasm and the like, but I don’t think it has those connotations here.

Φ�σις is apparently derived from φ�ω, which means, among other things, to grow or put forth shoots. I don’t think ‘blossom’ would be a bad translation in many instances. Another derivative is φυτ�ν, which means ‘plant’ as opposed to animal (ζω̃�ον).

While φ�σις does come to mean something like ‘fixed character’, I wonder (particularly in a NYC where spring is bounding outward) how much the word retains the conotations of growth.

Ο�σ�α is, of course, derived from the word for ‘to be’. But I suspect that in philosophical use φ�σις and ο�σ�α are used almost interchangeably.^{13}

Mr Pierce writes:

In the latter passage, Smith treats the φ�σις κα� ο�σ�α as a ‘hendiadys’, if that is the term, a one-thing-through-two-words: ‘essential nature’. Fair enough, I suppose; then Aristotle is making not a tripartite, but a bipartite division.

This is an interesting idea. I can’t tell whether this is a hendiadys or not. Anyone else have any thoughts? (I don’t think it’s a grammatical question, but one of the meaning of the terms.)

By the way, Hicks seems to think this a hendiadys. He glosses the phrase κα� τ�ν ο�σ�αν as ‘that is to say, its essence’, remarking that κα� is ‘explicative’. I don’t see what he then does with the τ� which appears earlier; I would have thought the τ� and κα� were linked.

91 (Mr Pierce). When I ordered a book called De Anima with ‘author’ W.D. Ross, I should have considered that it might be the Oxford Text, not Ross’s commentary on that text. Indeed, the Oxford Text is what I have now. Since we are moving slowly, I want to see that I understand the Greek of the first few sentences. Here is that Greek, with a translation into ‘Eek’—if that is what to call a nearly word-for-word rendering into English. I also refer to some paragraphs of Smyth’s Greek Grammar, and otherwise make some parenthetical comments. I’m just an amateur scholar of course, so I welcome corrections, as well as explanations of how to use, for example, Ross’s ‘apparatus criticus’. (The latter seems to be pretty clear, generally, but as I say below, I’m not sure how to read Ross’s very first note.)

I infer from Ross’s ‘Sigla’ (translation?) that his source is five manuscripts, in Paris and the Vatican and ‘Ambr.’ (where is that?), the oldest being from the ninth century, the youngest the 14th. So even the oldest text is closer to us than to Aristotle. (Mr Thomas mentioned the possibility of medieval intrusions into the texts that ought to be considered if one is really to understand the Philosopher himself.)

The first sentence, in three parts:

Τω̃ν καλω̃ν κα� τιμ�ων τ�ν ε�δησιν �πολαμβ�νοντες

Among the fine and honorable things the knowledge supposing (that is, ‘[While we] suppose [all] knowledge to be fine and honorable’)

μα̃λλον δ� �τ�ραν �τ�ρας � κατ� �κρ�βειαν � τω̃� βελτι�νων τε κα� θαυμασιωτ�ρων ε˜�ιναι

more some than other, either according to exactness, or by of-better-and-more-amazing being (Does Ross’s note mean that this phrase is omitted by some sources? My ‘according to’ is by Smyth 1676 and 1690. The τω̃� ε�ι̃ναι ‘by being’ seems to be an instrumental dative, specifically a ‘dative of standard of judgment’, Smyth 1512; but some texts have a genitive plural article τω̃ν instead, or none—I suppose Ross’s choice is correct.)

13Indeed, LSJ says as much.

δι� �μφ�τερα ταὺτ τ�ν περ� τη̃ς ψυχη̃ς �στορ�αν ε�λογως �ν �ν πρ�τοις τιθε�ημεν.

owing to both these, the about-the-soul account well among foremost might we place. (The verb is an optative with �ν, a ‘potential optative’, Smyth 1824.)

So: ‘Supposing that knowledge is one of the fine and honorable things,—some knowledge more than others, whether (a) because of its exactness or (b) because it is knowledge of better and more wonderful things,—for both reasons, we would do well to place an account of the soul in the forefront.’

Second sentence:

δοκει̃ δ� κα� πρ�ς �λ�θειαν �πασαν � γνω̃σις α�τη̃ς μεγ�λα συμβ�λλεσθαι, μ�λιστα δ� πρ�ς τ�ν φ�σιν·

Seems even to truth-all the knowledge of it greatly to contribute, greatest to the nature; (I note the formally parallel usage of ‘truth’ and ‘nature’, although the meanings may not be parallel. A contribution to truth is presumably a contribution to the stock of truths that we know; a contribution to nature would be a contribution to our stock of truths about nature. But I don’t know that ‘stock of truths’ is the best way to understand �λ�θεια �πασα.)

�στι γ�ρ ο�ι̃ον �ρχ� τω̃ν ζ�ν.

maybe, indeed, sort-of principle of the living. (The verb �στι, so accented, is ‘quasi-impersonal’, says Smyth 1985; for the LSJ lexicon, a mere ‘impersonal’ is enough; each reference-work gives a connotation of possibility.)

Perhaps then: ‘It seems that a knowledge of soul would contribute greatly to truth collectively, and especially concerning nature; for it may be that soul is a sort of principle of living things.’ Aristotle sure seems to avoid committing himself.

The third sentence:

�πιζητου̃μεν δ� θεωρη̃σαι κα� γνω̃ναι τ�ν τε φ�σιν α�τη̃ς κα� τ�ν ο�σ�αν,

We seek to consider and to know the nature of it and the essence, (The Greek and English word-orders coincide—except for the particle τε that Mr Thomas mentioned.)

ε˜�ιθ� �σα συμβ�βηκε περι α�τ�ν·

then as-much-as has-come-together near it;

˜ �ων τ� μ�ν �δια π�θη τη̃ς ψυχη̃ς ε˜�ιναι δοκει̃,

of which some peculiar accidents of the soul to-be seem,

τ� δ� δι� �κε�νην κα� τοι̃ς ζ�οις �π�ρχειν.

others owing-to that to-the-living to-belong.

‘We seek to contemplate and to know both its nature and essence; then, its accidents—of which (a) some seem to be peculiar to the soul, while (b) others seem to belong to living things because they have a soul.’

92 (Mr Thomas). Mr Pierce translates ζω̃�ον as ‘living thing’, but I think it would have been heard as ‘animal’. I have no access to the LSJ at the moment, but wonder whether there are occurances of the word in the more general sense prior to Aristotle.

93 (Mr Pierce). If I understand Mr Thomas about ζω̃�ον, he suspects that the word before Aristotle meant ‘animal’, but Aristotle allowed it to refer to living things generally, including plants. I hadn’t considered such a possibility. The LSJ lexicon (available at http://www.perseus.tufts.edu/) gives the definition ‘living being, animal’, but doesn’t seem to give any instance that refers to plants (unless I haven’t studied carefully enough). Indeed, a usage of Plato’s is cited that distinguishes ζω̃�α from φυτ�.

But also, LSJ give an example from Herodotus wherein ζω̃�α refers to pictures of the bridging of the Bosphorus.

94 (Mr Schneider). From the plane, no less, Mr Thomas muses:

Mr Pierce translates ζω̃�ον as ‘living thing’, but I think it would have been heard as ‘animal’. I have no access to the LSJ at the moment, but wonder whether there are occurances of the word in the more general sense prior to Aristotle.

There’s a similar question about the uses of the Hebrew word cha-yot, which might be literally translated ’life-things,’ but is often seen as ’creatures.’ Might that do here?

95 (Mr Salas). §1. Mr Pierce had commented [in (93)] on the use of the noun and adjective formed from the verb ζωαω[.] This has been on my mind since yesterday and I have a theory that may or may not pan out.

First, the stuff that’s useful and not speculative. I think that Herodotus, and others, use ζ�α, the neuter plural of the noun in paintings as metaphorical or virtual living things. I’m not certain about this one but I don’t think you could have this use with a painting of a totally inanimate object. Does our word portrait function similarly? I’ll look around some more.

Now for some speculation! I’ve found three words that look nearly alike in Greek and all mean slightly different things. They are the noun ζω̃�ον, the adjective ζω�ς, -�, -�ν, and the participle ζ�ων (and its other forms). The only difference between the noun and the adjective is that the accent of the noun is on the penult while the accent of the adjective is on the ultima. So, in the text, one can tell when the manuscripts prefer one reading over another.

Finally, depending on what form the participle takes, it can appear as either the adjective or the noun. The reason this may be important is that participle and the adjective most likely are consistent in meaning ‘living x’. I’m hoping to find that Aristotle consistently uses one of these two terms when he’s talking about the broad class of living things, τ� ζ�οντα. Note that that is the participle functioning substantively. In the case of the noun, I suspect that the text will be less careful (if it’s being careful at all) but I hope that any instances of contrast between plants and animals will be between τ� φυτ� and τ� ζω̃�α, the noun.

If these distinctions hold, we might have a little more clarity about which sort of beings we’re talking.

Mr Pierce, I’ll take a look tomorrow but if you want to look as well, take a look at the uses of the adjective and the verb on the LSJ and see if they give elucidate our puzzle. I will also look at Ross.

§ 2. [Regarding (91):] Some thing that may prove useful to some of you since Mr Pierce has mentioned it. If you don’t own it, Smyth is on-line and accessible at Perseus.

‘Ambr.’ stands for Ambrosianus and it’s dated from 12–13 a.d. I don’t know where it is, since it doesn’t have a place name like Vaticanensis or Parisensis or some such. y, or Par. 2034 might date from the 8th century.

In his commentary, Ross cites 11 manuscripts but only five major families. E, L, and P all seem to belong to different families while SUVx and CWy seem to spring from one archetype but form two different sub-families.

Check this link out, it looks neat and informative but does not give a full list of

apparatus criticus abbreviations: http://www.brocku.ca/classics/parker/
LATI3V14/Apparatus_Criticus.pdf

However, the pdf does mention that the big LSJ has a list of abbreviations in the front of it.

Testimonia and ancient commentaries may alleviate this problem [of ‘medieval intrusions into the texts] but I don’t see any way of really gettting around it.

The greater manuscript support for the dative may be the reason that Ross retains the dative here, although it confused me at first glance. Also, Alexander omitted the line from μα̃λλον to ˜�ειναι but that’s contained in scholia/lemmata/ commentary(?) but for no textual reasons. The manuscripts don’t see[m] to have a problem with it.

[Concerning the first sentence:] I don’t know whether your ‘would’ translated the potential optative. I thought that ‘could’ would be better. I would translate, ‘On account of these things, we could reasonably place [my/the] account of the soul in the forefront.’

[Concerning the second sentence:] I disagree with the way in which you’re reading ο˜ �ιον. In this sentence, ο˜ �ιον can’t be modifying �ρχ� because �ρχ� is nominative feminine singular and ο˜ �ιον is either masculine accusative singular, neuter accusative singular, or more likely nominative neuter singular. ο˜ �ιον �στι can often idiomatically mean ‘for example.’ It’s a use especially common in Aristotle. So, I think the sentence should read, ‘For example, because it is the principle/fount of living things.’ My reading doesn’t give Aristotle that much wiggle room.

I have to get to sleep but I’ll write more later, especially about the question of hendiadys in the upcoming sentence.

96 (Mr Pierce). Mr Salas mentioned that the Smyth grammar is at Perseus. It is also available in a different format at http://www.ccel.org/s/smythe/ grammar/html/toc_uni.htm. I don’t think I could do without the tangible volume I have at home though. I’m wondering whether to splurge on the big LSJ when we go to istanbul in a couple of days—it’s available there. The intermediate lexicon that we used at St John’s doesn’t have the word ε�δησιν from the very first line of text. That the word means ‘knowledge’ is perhaps obvious, but still!

Thanks, Mr Salas, for the link about apparatus critici; it was useful. A websearch suggests that ‘Ambrosianus’ refers to the Ambrosian Library in Milan.

You suggest, I think, that if Aristotle wants to distinguish between animals and living things in general, he can use the participle ζωοντα for the latter, and the noun ζ�α for the former. Perseus is offline while most Americans are sleeping, so I can’t check the LSJ. I should have recognized before that already in (90) Mr Thomas had mentioned the distinction between ζωα and φυτα, the latter word, for plant, being related to φ�σις.

Concerning my translation of the first sentence of De Anima, you [are] certainly right that ‘would’ is wrong. I think I meant to say ‘might’, where you propose ‘could’; ‘might do well’ sounds more natural to me than ‘could do well’, but I don’t suppose it’s an important distinction.

About the end of the second sentence: I was just guessing that ο�ον had become an adverb meaning ‘sort-of’, that is, ‘roughly’ or ‘approximately’. I had been noticing what I imagined was a free and elliptical style in Aristotle, exemplified in the upcoming λεγω. This is the ‘I-say’ of Euclid, but for Aristotle it seems to be ‘I mean’ or just ‘i.e.’

Now Mr Salas, you give Aristotle’s ‘for-example’ construction as ο�ον εστι; but the text at the point in question has the opposite order �στι ο�ον, the �στι having the unusual accentuation which the scholars say connotes possibility. Do you think such a connotation is not present here?

97 (Mr Thomas). The Greek text that Mr Pierce acquired was Ross’s OCT, published in 1956. The text that Mr Salas and I are using is the text in Ross’s edition and commentary from 1961. The apparatus is different in the two. I don’t know if there are any textual differences.